OpenStudy (anonymous):

Write the percent comparing the new amount to the original amount. Then find the new amount. 75 feet increased by 4%. I know the new amount is 105 ft but i don't understand the directions :/

OpenStudy (anonymous):

*new amount is 78 ft Sorry. Unless I'm mistaken. Please correct me and explain the answer thank you!

OpenStudy (anonymous):

yes it is 78 ft. .04*75=3 3+75=78. the percent should be 104% i think 104/100=78/75

OpenStudy (anonymous):

Thank you! I have one more only if you want to help. It's 48 hours decreased by 16%.